OpenStudy (anonymous):

The work of a student to find the dimensions of a rectangle of area 8 + 16x and width 4 is shown below: Step 1: 8 + 16x Step 2: 4(4) + 4(12x) Step 3: 4(4 + 12x) Step 4: Dimensions of the rectangle are 4 and 4 + 12x In which step did the student first make an error and what is the correct step? Step 3; 4 + (4 + 12x) Step 3; 4 + (4 ⋅ 12x) Step 2; 4(2) + 4(4x) Step 2; 4(2) + 4x(2)

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@DecentNabeel

OpenStudy (calculusxy):

To find the area of a rectangle what you do is: length x width = area You already have the width as 4 and area: length x 4 = 8 + 16x \[l \times 4 = 8 + 16x\] To solve for l (length) you need to isolate and use the division property of equality: \[\frac{ l \times 4 }{ 4 } = \frac{ 8 + 16x }{ 4 }\] Clearly the student did the inverse in one of the steps. Can you find out which step he/she did that?

OpenStudy (anonymous):

C Step 2; 4(2) + 4(4x)

OpenStudy (calculusxy):

Yes

OpenStudy (anonymous):

Yay c:

OpenStudy (calculusxy):

Great job!

OpenStudy (anonymous):

Thanks for your help (: